OpenStudy (emilyjones284):

Evaluate the integral. (8cosx+6sinx-x^2/2)dx

OpenStudy (anonymous):

\[\int\left( 8\cos(x)+6\sin(x)-\frac{x^2}{2}\right)dx\]?

OpenStudy (emilyjones284):

Yes